OpenStudy (anonymous):

find the slope and the y-intercept of each line whose equation given -4x+y=7

OpenStudy (anonymous):

y = 7+ 4x y = 4x + 7 now its in the form of y = kx +d where k is the slope slope = 4

OpenStudy (precal):

y intercept is 7